OpenStudy (anonymous):

Simplify 4 radical to the 3rd (135)

OpenStudy (saifoo.khan):

\[\huge 4\sqrt[3]{135}\]

OpenStudy (saifoo.khan):

this?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

\[\sqrt[4]{135^3}= \sqrt[4]{(3^3\times5)^3}= \sqrt[4]{3^6\times 5^2}=3\sqrt[4]{3^2\times 5^2}\]

OpenStudy (lalaly):

\[12\sqrt[3]{5}\]

OpenStudy (anonymous):

thxs!

OpenStudy (saifoo.khan):

Lala's right.